OpenStudy (anonymous):

find root for the following: f(x)=x+1/x−1

OpenStudy (anonymous):

to find the root, put f(x)=0 so, x+1/x−1 = 0 x=.... ?

OpenStudy (anonymous):

is it the correct: f(x)=x+1/x-1 =0, x+1=0 x=-1 ???????????

OpenStudy (anonymous):

yes, absolutely correct :)

OpenStudy (anonymous):

thanks for confirming

OpenStudy (anonymous):

welcome ^_^